# Multi-label Scene Classification for Autonomous Vehicles: Acquiring and Accumulating Knowledge from Diverse Datasets

### v2

@@ Abstract
Driving scene identification, which assigns multiple non-exclusive class labels to a scene, provides the contextual awareness necessary for enhancing autonomous vehicles’ ability to understand, reason about, and interact with the complex driving environment. As a multi-label classification problem, it is better tackled via multitasking learning. However, directly training a multi-label classification model for driving scene identification through multitask learning presents two main challenges: acquiring a balanced, comprehensively annotated multi-label dataset and balancing learning across different tasks. This paper introduces a novel learning system that synergizes knowledge acquisition and accumulation (KAA) with consistency-based active learning (CAL) to address those challenges. KAA acquires and accumulates knowledge about scene identification from various single-label datasets via monotask learning. Subsequently, CAL effectively resolves the knowledge gap caused by the discrepancy between single-label and multi-label data. An ablation study on our Driving Scene Identification (DSI) dataset demonstrates a 56.1% performance increase over the baseline model pretrained on ImageNet. Of this, KAA accounts for 31.3% of the gain, and CAL contributes 24.8%. Moreover, KAA-CAL stands out as the best performer when compared to state-of-the-art (SOTA) multi-label models on two public datasets, BDD100K and HSD, achieving this while using 85% less data. The DSI dataset and the implementation code for KAA-CAL are available at https://github.com/KELISBU/KAA-CAL.
Index Terms:
Autonomous Vehicle, Driving Scene Identification, Multi-label Classification, Knowledge Distillation, Active Learning
@@ I Introduction
Driving scene identification involves assigning one or multiple non-exclusive class labels to a scene captured by an onboard camera, selecting from a pre-defined set of categories. For example, driving scenes can be identified based on the weather condition, with possible exclusive categories including clear, overcast, cloudy, raining, and snowing. Scene identification is a crucial function for autonomous vehicles(AVs), as it provides the context or situational awareness for decision-making [1, 2, 3]. For example, the friction coefficient of the road surface varies, depending on whether the road is dry, wet, snowy, or icy. Identifying driving scenes based on the road surface condition allows vehicles to adjust the estimations of braking distance and turning speed accordingly, thus improving safety. Scene identification also provides strong support for complex scene reasoning and understanding. From a statistical perspective, contextual awareness enables the estimation of conditional probabilities for random events that influence driving decisions. Conditional probabilities provide more accurate assessments than overall probabilities when the identified conditions are relevant to the event. For example, the probability of a deer suddenly appearing on the road is generally low. However, this probability increases significantly at dawn or dusk, particularly on two-way roads in rural areas. By identifying these conditions, an animal detection system can adaptively adjust its decision threshold, enhancing its ability to respond cautiously.
Driving scene identification is an image classification problem. Due to their complexity, driving scenes can be classified according to multiple scene attributes, resulting in a multi-label classification problem. Each scene attribute is a categorical variable with mutually exclusive yet collectively inclusive classes. Example attributes include weather, time of day, weather-related road condition, road function system, among others. This multi-label image classification problem can be addressed either through monotask or multitask learning. In the monotask learning approach, multiple classification models are trained independently, each classifying driving scenes based on one attribute and thus assigning one class label to an input image. In contrast, multitask learning integrates all the classification tasks into one single model, enabling the simultaneous assignment of multiple non-exclusive class labels to an input image. A multitask model is generally preferred over a set of monotask models because the multitask model is more computationally efficient and, sometimes, can leverage task correlations to improve learning [4, 5].
@@
However, training a multitask model for complex driving scene identification is way more difficult than training a set of monotask models for several reasons. A primary challenge is the lack of a comprehensively annotated multi-label training dataset. In the high-dimensional attribute space, driving scene distribution is highly imbalanced [6, 7]. Some categorical combinations are rare, making it difficult to collect training samples. This data scarcity can cause the model to perform less reliably in low-frequent scenarios. Additionally, a multitask model may struggle to balance learning across tasks, especially when competing gradients arise. These challenges likely explain why image- and video-level annotations in popular driving scene datasets are either single-labeled (e.g., [8, 9]) or multi-labeled with limited scene attributes (e.g., [10, 11]), and why existing models are predominantly single-label classification models (e.g., [12, 13, 14]).
The challenge in collecting a balanced multi-label training dataset for complex driving scene identification, combined with the abundance of single-label data sources, motivates the exploration of a novel idea: acquiring knowledge about complex scene identification through monotask learning across diverse single-label datasets. Existing approaches using single-label classification to enhance multi-label scene classification still rely on a multi-label training dataset [6, 7]. Consequently, they are not applicable to the research problem in this paper, which focuses on learning across different datasets. The approach that this paper is interested in is analogous to the learning process of students who study various subjects by adhering to a recurring weekly schedule throughout an academic year. The central idea of the learning process is as follows. A student model learns one classification task from a corresponding single-label dataset at a time in a sequential manner and accumulates the acquired knowledge in a teacher model that serves as a reference for the student model’s continual learning and improvement. This process repeats cyclically until the student model’s performance saturates.
However, learning multi-label classification from single-label datasets through monotask learning must address a domain shift problem, stemming from the discrepancy between the marginal distributions of individual attributes and their joint distribution. This issue could be potentially mitigated by developing an active learning algorithm that selects and annotates some representative samples from the high-dimensional space of scene attributes. Those samples can teach the trained classification model to adapt.
@@
  • A new deep learning system, Knowledge Acquisition and Accumulation (KAA), is proposed for complex driving scene identification.
  • In demonstrating KAA, a comprehensive Driving Scene Identification (DSI) dataset is constructed, comprising seven single-label subsets, each annotated according to one specific scene attribute.
  • An adaptation algorithm, Consistency-based Active Learning (CAL), is developed to address the domain shift problem confronted by KAA.
Remainder of the paper is organized as follows. Section II summarizes recent work related to this study. Then, Section III details the components of the proposed DSI dataset. After that, the learning framework KAA and the adaptation algorithm CAL are introduced in Section IV, with the implementation delineated in Section V. Section VI further reports the experimentation and results. At the end, Section VII concludes the study by summarizing important findings and future research directions.
@@ II Related Work
Existing work that builds a foundation for this study includes public driving scene datasets, multi-label scene classification from a computer vision perspective, knowledge distillation, and deep active learning. The most relevant literature is discussed below.
@@ II-A Public Driving Scene Datasets
Various driving video datasets have been developed to support different perception tasks, including scene classification. BDD100K [10] is a large-scale driving video dataset created by the Berkeley DeepDrive (BDD) research group, featuring image-level annotations for six weather conditions, six scene types, and three distinct times of day. Honda Scene Dataset (HSD) [11] contains 80 hours of diverse, high-resolution driving video data clips collected in the San Francisco Bay Area. HSD provides video-level identification of eleven road places, four categories of road environment, and four weather conditions, thereby broadening the diversity of both attributes and classes for scene identification. ROADWork [9] is a large-scale public dataset focused on the identification of work zones with data collected from eighteen cities in the U.S. This dataset also provides fine-grained annotations for instance and semantic segmentation. DENSE++ [4] provides multi-label annotations for environmental conditions including daytime, precipitation, fog, road condition, roadside condition, and scene-setting for 12,997 images collected from Northern Europe. The annotation method is provided, but the labels are not publicly available. WZ-Traffic [8] is a scene dataset with 6,035 single-label images in 20 categories. Other open-source driving datasets, including Cityscapes [15], KITTI [16], and nuScenes [17], are primarily annotated for object detection, semantic segmentation, and instance segmentation tasks.
A review of those datasets reveals that few provide comprehensive annotations for multi-label scene classification. Some datasets provide class labels either at the frame level or the video level, yet they do so based on one or a limited number of attributes of their interest. Additionally, there is no common standard or widely accepted taxonomy for categorizing driving scenes with respect to any given scene attribute, leading to inconsistencies in scene classes across different datasets.
@@ II-B Multi-label Driving Scene Classification
A few studies have tackled the problem of multi-label scene classification in various approaches. Duong et al. [6] developed CF-Net, which uses single-label classifiers to enhance the main multi-label classifier by utilizing feature fusion and stacking. The model was trained on a modified BDD100K dataset with three attributes: location, weather, and time of day. Chen et al. [7] proposed to solve the multi-label scene classification problem by incorporating the single-label training procedure into the multi-label architecture. That allows for addressing the issue of imbalanced categories. Additionally, a deep data integration strategy was utilized to improve the classification ability. RECNet [4] was proposed as a multi-label classification model for identifying environmental conditions. It is a multitask model that uses the EfficientNet-B2 as the shared backbone and six downstream classifiers to identify all annotated environmental conditions, including daytime, fog, precipitation, road condition, roadside condition, and infrastructure. The study proposed a hierarchical strategy for annotating the environmental conditions, generating a fully annotated dataset DENSE++ for training the multi-label classification model. Prykhodchenko and Skruch [5] also trained a multi-label scene classification model on BDD100K via multitask learning. Data augmentation and balancing strategies were utilized to address the class imbalance issue in the BDD100K dataset.
A review of those studies on multi-label scene classification suggests that multitask learning is a more commonly pursued approach than monotask learning for training such models. Integrating knowledge from single-label classification into multi-label classification models has been shown to be an effective strategy for performance improvement. To the best of our knowledge, no multi-label classification models have been developed through monotask learning from diverse datasets.
@@ II-C Knowledge Distillation
Knowledge distillation (KD) typically utilizes a teacher-student framework to distill the knowledge accumulated in a deep or large model (the teacher) into a smaller or shallow one (the student). It can support the student network’s continual acquisition of scene identification knowledge.
KD methods are generally categorized into three types: response-based KD, where the student directly mimics the teacher’s final predictions; feature-based KD, which transfers knowledge by leveraging features extracted by the teacher; and relation-based KD, which explores relationships between different layers or samples [18]. When both the teacher and student networks are deep, Romero et al. [19] demonstrated that regulating multiple hidden layers of the student network leads to superior performance compared to response-based KD. Ma et al. [20] introduced ARK, a novel framework that accrues and reuses feature-based knowledge to address the challenge of inconsistent annotations in various medical image datasets. Using ARK, a foundation model is constructed for medical image analysis. However, when KD occurs progressively, there is a risk of forgetting previously acquired knowledge. To mitigate this issue, Van de Ven et al. [21] proposed a brain-inspired method that replays internal or hidden representations in continual learning. While sharing some similarities with proposed KAA, ARK primarily focuses on addressing annotation heterogeneity across different datasets.
@@ II-D Deep Active Learning
Deep active learning is typically designed to improve labeling efficiency when faced with the high cost of annotation. Sample selection is guided by an uncertainty-based, diversity-based, or hybrid method.
An uncertainty-based method employs a query strategy to select samples, labeling uncertain samples that provide the most informative data points. Methods such as maximum entropy, margin sampling, least-confidence sampling, and Bayesian Active Learning by Disagreement (BALD) rely on information derived from a single task [22]. Additionally, a task-agnostic method introduced in [23] integrates a loss prediction module to select top samples of the highest prediction loss within the unlabeled pool, prioritizing them for further downstream tasks.
A diversity-based method selects instances from the unlabeled pool to represent a broad and varied data distribution. Beyond traditional clustering methods like K-means, Core-Set Section [24] computes the distance of features from a designated hint layer within a deep learning model. A greedy search algorithm is employed to iteratively select samples that exhibit the greatest distance from their nearest neighbors until the selection budget is exhausted.
Hybrid methods, including Wasserstein Adversarial Active Learning (WAAL) and Batch Active learning by Diverse Gradient Embeddings (BADGE), effectively mitigate sample redundancy caused by uncertainty-based methods. They also overcome the limitation of diversity-based methods, which often overlook the most informative instances [22]. Hekimoglu et al. [25] introduced a hybrid method for multitask learning, where uncertainty is quantified by an inconsistency score computed as the maximum loss between the initial and refined pairwise task predictions. Simultaneously, a reconstructed feature embedding that condenses information across all tasks is employed to measure the diversity score. The combination of these two components forms the deep active learning basis for multitask learning. However, this approach is limited to multi-label datasets and does not address the domain shift in the high-dimensional feature space that arises when merging multiple single-label datasets.
@@ III Driving Scene Identification (DSI) Dataset
The data distribution in the high-dimensional scene attribute space is highly imbalanced. Consequently, some combinations of scene classes are corner cases and challenging to collect data for. This limits the creation of comprehensively annotated datasets to directly support multi-label scene classification. However, collecting data independently from the marginal distributions of scene attributes is relatively easier than sampling from their joint distribution.
This paper introduces DSI, a single-label dataset created to demonstrate the proposed learning systems KAA-CAL. DSI consists of 31,835 scene images sampled from public driving video datasets, including BDD100K [10], HSD [11], and ROADWork Data [9], and is supplemented with additional images from YouTube videos. An overview of the DSI dataset and its statistics are presented in Fig. 1. The dataset is available to the public at https://github.com/KELISBU/KAA-CAL.
@@ III-A Driving Scene Attributes and Classes
The DSI dataset consists of $M$ (7) subsets, denoted as $\{\mathcal{D}_{m}\}_{m=1}^{M}$. Each subset is annotated according to one unique scene attribute and further partitioned into training (Trn), validation (Vld), and test (Tst) sets:
$\mathcal{D}_{m}=\mathcal{D}_{m}^{Trn}\cup\mathcal{D}_{m}^{Vld}\cup\mathcal{D}_{m}^{Tst}.$ (1)
where $m$ is the index of scene attributes.
@@
As illustrated in Fig. 1, these scene attributes include Grade-separated Infrastructure, Road Function, Weather, Work Zone, Weather-related Road Condition, Time of Day, and Intersection Type. In total, DSI provides image-level labels for 24 classes, with each falling exclusively under one of the 7 scene attributes. Classifying driving scenes according to these scene identification attributes offers valuable insights for AVs.
@@
Road segments featuring grade-separated infrastructure, such as bridges and tunnels, often present specific challenges for AVs. Bridges above the road may limit the sensor field of view, cause sudden lighting changes, and introduce gusts of wind. Tunnels, on the other hand, have different lighting conditions, limited visibility, and potential congestion. DSI includes a Grade-separated Infrastructure subset comprising Overhead bridges, Tunnels, and Open roads, totaling 7,765 images. This subset was curated from the HSD dataset using query keywords like “overhead bridge” and “tunnel”. To ensure these structures are visually perceivable, additional query criteria such as “Approaching” and “Entering” were used for data sampling.
@@
Each road type has specific characteristics, such as a speed limit range, traffic density, vehicle type distribution, terrain, and travel purposes. Thus, understanding the road function assists adaptive driving behavior. The US road functional system has not been systematically annotated in public datasets. To address this, the Road Function subset was created by querying with keywords like “road function” on YouTube and sampling every fifth frame from the collected videos. This dataset consists of 5,740 images across 4 classes: Local, Arterial, Interstate, and Collector.
@@
Identification of the weather condition is critical because an adverse weather condition risks driving safety and the performance of visual tasks. The Weather subset includes 4,698 frames sampled from the BDD100K dataset, encompassing Overcast, Clear, Foggy, Snowing, and Raining conditions. To mitigate the issue of data imbalance caused by the rarity of foggy scenes, synthetic images were generated to ensure a more balanced distribution across weather conditions.
@@
Driving safety is a concern in work zones, where construction equipment, temporary barriers, road workers, and other related facilities are present. Work zones often involve lane closures. Confusion arises when old, not fully removed lane markings mix with new ones, especially for AVs. Identifying work zones allows AVs to adjust their lane position and speed dynamically, minimizing crash risks. The training and validation data of the Work Zone subset mainly come from BDD100K and HSD, and its testing data are primarily from the ROADWork dataset.
@@
The friction coefficient of the road surface changes depending on weather conditions like snow or rain. Identifying these weather-related road conditions allows AVs to adjust their control for better adaptation to the specific surface characteristics. The Weather-related Road Condition subset includes 3,693 images collected from both YouTube and BDD100K, intended for identifying Snowy, Dry, and Wet conditions.
@@
The lighting condition directly impacts the visibility of objects, roads, and the surrounding environment because illumination influences key visual functions like contrast sensitivity, visual acuity, depth perception, and peripheral vision. Consequently, the lighting condition affects human drivers’ visual perception and reaction times. It also impacts machine vision due to the affected quality of camera-captured data. For example, nighttime and dawn/dusk images often exhibit increased noise, reduced contrast, and color distortion. The Time of Day subset comprises 2,978 images labeled as Night, Daytime, or Dawn/Dusk, sampled from the BDD100K dataset.
@@
More than half of the combined total of fatal and injury crashes occur at or near intersections due to increased conflict points and complex interactions among vehicles, pedestrians, and other road users. Recognizing the type of intersection allows AVs to anticipate potential conflict points and make informed decisions, such as yielding, stopping, or merging in complex environments. The Intersection Type subset consists of 2,680 images, categorized into four classes: None, 3-way, 4-way, and Roundabout. The data were collected by integrating YouTube videos, which provide roundabout images, along with the HSD and BDD100K datasets, which contribute images of none, 3-way, and 4-way intersections. 3-way and 4-way intersections are typically more noticeable when ego-vehicles are entering the intersection, but less so when approaching or passing through them. Therefore, the query keyword “entering” was used when searching for images of 3-way and 4-way intersections in the HSD dataset.
III-B Diversity and Generalization
Small biases in training data can lead to poor generalization of deep learning models, as a biased dataset may cause the model to learn features that deviate from the intended task [26]. Therefore, the data collection heavily emphasized diversity and generalization to mitigate potential biases. For instance, in the Work Zone subset, data points were sampled across various land use categories (urban and rural areas), road functions, and lighting conditions. This ensures that the DSI dataset provides a good opportunity for learning representations likely encountered in the real world.

### v3

@@ Abstract
Driving scenes are inherently heterogeneous and dynamic. Multi-attribute scene identification, as a high-level visual perception capability, provides autonomous vehicles (AVs) with essential contextual awareness to understand, reason through, and interact with complex driving environments. Although scene identification is best modeled as a multi-label classification problem via multitask learning, it faces two major challenges: the difficulty of acquiring balanced, comprehensively annotated datasets and the need to re-annotate all training data when new attributes emerge. To address these challenges, this paper introduces a novel deep learning method that integrates Knowledge Acquisition and Accumulation (KAA) with Consistency-based Active Learning (CAL). KAA leverages monotask learning on heterogeneous single-label datasets to build a knowledge foundation, while CAL bridges the gap between single- and multi-label data, adapting the foundation model for multi-label scene classification. An ablation study on the newly developed Driving Scene Identification (DSI) dataset demonstrates a 56.1% improvement over an ImageNet-pretrained baseline. Moreover, KAA-CAL outperforms state-of-the-art multi-label classification methods on the BDD100K and HSD datasets, achieving this with 85% less data and even recognizing attributes unseen during foundation model training. The DSI dataset and KAA-CAL implementation code are publicly available at https://github.com/KELISBU/KAA-CAL.
@@ I Introduction
Driving scene identification involves assigning non-exclusive labels to each scene captured by an onboard camera. As illustrated in Fig. 1, a complex scene is characterized using various attributes, such as time of day, weather, roadway function, intersection type, among others. Scene identification is a high-level perception crucial for autonomous vehicles (AVs), as it provides the contextual awareness for reasoning and decision-making [1, 2]. For example, the friction coefficient of the road surface varies depending on whether the road is dry, wet, snowy, or icy. Identifying driving scenes based on the road surface condition allows vehicles to adjust their estimations of braking distance and turning speed accordingly, thus improving safety. Driving scene identification also supports data generation, as it provides scene attribute labels that serve as semantic priors to guide image and video synthesis. For example, after aligning visual and semantic representations, the intersection type class serves as an explicit conditioning information for fine-grained generation (e.g., three-way, four-way, or roundabout).
Figure 1: A foundation model for multi-attribute driving scene identification, with its comprehensive knowledge obtained from diverse single-label datasets
Driving scenes are inherently heterogeneous and dynamic, requiring the simultaneous recognition of many attributes. This makes scene identification a multi-label classification problem that goes far beyond a simple perception task. The complexity nature of this important problem highlights the critical need for a foundation model possessing the comprehensive knowledge to accurately characterize these intricate scenes.
@@
While seemingly straightforward for driving scene identification, standard multitask learning methods face critical limitations. A primary challenge is the lack of a comprehensively annotated multi-label training dataset. In a high-dimensional attribute space, driving scene distribution is highly imbalanced [3, 4]. Some scenarios are rare, making it difficult to collect balanced multi-label training samples. This likely explains why image- and video-level annotations in popular driving scene datasets are either single-labeled (e.g., [5, 6]) or multi-labeled with just limited scene attributes (e.g., [7, 8]), and why existing models are predominantly single-label classification models (e.g., [9, 10, 11]). Meanwhile, as the transportation system evolves, the need to identify additional attributes constantly emerges. However, the standard multitask learning framework proves inefficient for integrating desired new knowledge. This is because it requires not only re-labeling entire training datasets with the required new labels but also a further data collection, a process that is neither scalable nor cost-effective.
The aforementioned challenges, combined with the abundance of single-label datasets and the convenience of monotask learning, motivate the exploration of a learning approach leading to a driving scene foundation model. As illustrated in Fig. 1, this strategy constructs the model by learning from heterogeneous datasets and consolidating transferable representations. The resulting model can not only recognize complex driving scenes using a variety of attributes but also quickly learn to recognize new ones.
The learning process to be introduced in this paper is inspired by how students learn. A student model concentrates on learning one classification task from a corresponding single-label dataset at a time in a sequential manner and consolidates the acquired knowledge in a teacher model, which serves as a reference for the student model’s continual learning and improvement. This sequential process goes cyclically, progressively transforming knowledge from heterogeneous single-label datasets into a unified representation, forming a foundation model for multi-attribute driving scene identification. Yet, this approach confronts an inevitable domain shift problem, which arises from the discrepancy between the marginal distributions of individual attributes and their joint distribution. To mitigate this, a strategy is developed to guide the model’s adaptation, transforming the knowledge it acquires from diverse single-label datasets into a comprehensive understanding suitable for complex scene recognition.
@@
  • A new deep learning system, Knowledge Acquisition and Accumulation (KAA), is proposed to learn from heterogeneous single-label datasets and consolidate learned knowledge into a unified representation, laying the groundwork for multi-label driving scene identification.
  • In demonstrating KAA, a comprehensive and scalable, Driving Scene Identification (DSI) dataset is constructed, comprising seven single-label subsets, each annotated according to one specific scene attribute, representing heterogeneous sources of information about driving scenes.
  • An adaptation algorithm, Consistency-based Active Learning (CAL), is developed to couple with KAA, addressing the domain shift problem it confronts.
Remainder of the paper is organized as follows. Section II summarizes recent work related to this study. Then, Section III details the components of the new DSI dataset. After that, the proposed KAA-CAL learning method is introduced in Section IV, with the implementation delineated in Section V. Section VI further reports the experimentation and results. At the end, Section VII concludes the study by summarizing important findings and future research directions.
@@ II Related Work
This paper is built upon a spectrum of research areas. The most relevant literature from these fields is summarized below.
@@ II-A Public Driving Scene Datasets
Various driving video datasets have been developed to support different perception tasks for AVs, including scene classification. BDD100K [7] is a large-scale driving video dataset featuring image-level annotations for six weather conditions, six scene types, and three distinct times of day. Honda Scene Dataset (HSD) [8] contains 80 hours of driving video data clips collected in the San Francisco Bay Area. HSD provides video-level identification of eleven road places, four categories of road environment, and four weather conditions, thereby broadening the diversity of both attributes and classes for urban scene identification. ROADWork [6] is a public dataset focused on work zones identification with data collected from eighteen cities in the U.S. This dataset also provides fine-grained annotations for instance and semantic segmentation. DENSE++ [12] provides multi-label annotations for environmental conditions including daytime, precipitation, fog, road condition, roadside condition, and scene-setting for 12,997 images collected from Northern Europe. While the annotation method is provided, the image labels are not publicly available. WZ-Traffic [5] is a scene dataset with 6,035 single-label images in 20 categories. Other open-source driving datasets, including Cityscapes [13], KITTI [14], and nuScenes [15], are primarily annotated for object detection, semantic segmentation, and instance segmentation tasks.
Few datasets provide comprehensive annotations for multi-label scene classification. Some provide class labels either at the frame level or the video level, yet they do so based on one or a few attributes of their interest. Integrating the information in heterogeneous datasets as a unified representation of driving scenes is greatly desired.
@@ II-B Multi-label Driving Scene Classification
Multi-attribute scene identification is commonly modeled as a multi-label image classification proble,. A few studies have tackled this problem in various approaches. Duong et al. [3] developed CF-Net, which uses single-label classifiers to enhance the main multi-label classifier by utilizing feature fusion and stacking. The model was trained on a modified BDD100K dataset with three attributes: location, weather, and time of day. Chen et al. [4] proposed to solve the multi-label scene classification problem by incorporating the single-label training procedure into the multi-label architecture. Additionally, a deep data integration strategy was utilized to improve the classification ability. RECNet [12] proposed a hierarchical strategy for annotating six environmental conditions, including daytime, fog, precipitation, road condition, roadside condition, and infrastructure, leading to a fully annotated dataset DENSE++. This dataset was used to train a multi-label classification model, which uses the EfficientNet-B2 as the shared backbone for the six downstream classifiers. Prykhodchenko and Skruch [16] also trained a multi-label scene classification model on BDD100K via multitask learning. Data augmentation and balancing strategies were utilized to address the class imbalance issue in the BDD100K dataset.
Existing methods generally adhere to the standard multitask learning framework, thus relying on multi-label training datasets. New methodologies that can transcend this framework will offer an opportunity to utilize the abundant driving scene images or video data.
@@ II-C Knowledge Distillation
Knowledge distillation (KD) typically utilizes a teacher-student framework to distill the knowledge accumulated in a deep or large model (the teacher) into a smaller or shallow one (the student). It supports the student network’s continual acquisition and consolidation of scene identification knowledge.
KD methods are generally categorized into three types: response-based KD, where the student directly mimics the teacher’s final predictions; feature-based KD, which transfers knowledge by leveraging features extracted by the teacher; and relation-based KD, which explores relationships between different layers or samples [17]. When both the teacher and student networks are deep, regulating multiple hidden layers of the student network leads to improved performance compared to response-based KD [18]. Ma et al. [19] introduced ARK to accrue and reuse feature-based knowledge, addressing the challenge of inconsistent annotations in various medical image datasets. Using ARK, a foundation model is constructed for medical image analysis. When KD occurs progressively, there is a risk of forgetting previously acquired knowledge. To mitigate this issue, Van de Ven et al. [20] proposed a brain-inspired method that replays internal or hidden representations in continual learning. While sharing some similarities with proposed KAA, ARK primarily focuses on addressing annotation heterogeneity across different medical datasets.
@@ II-D Deep Active Learning
Deep active learning can efficiently adapt models to new domains or tasks, achieved by strategically selecting the most informative and representative samples for adaptation. Some uncertainty-based methods such as maximum entropy, margin sampling, least-confidence sampling, and Bayesian Active Learning by Disagreement rely on uncertainty measures based on a task [21]. In contrast, a task-agnostic method introduced in [22] integrates a loss prediction module to select top samples of the highest prediction loss from the unlabeled pool. Uncertainty-based methods may cause sample redundancy.
A diversity-based method selects instances from the unlabeled pool to represent a broad data distribution. Beyond traditional clustering methods like K-means, Core-Set Section [23] computes the distance of features from a designated hint layer within a deep learning model. A greedy search algorithm is employed to iteratively select samples that exhibit the greatest distance from their nearest neighbors until the selection budget is exhausted. Diversity-based methods may overlook the most informative instances.
Hybrid methods, including Wasserstein Adversarial Active Learning and Batch Active learning by Diverse Gradient Embeddings, effectively address limitations of both uncertainty- and diversity-based methods [21]. For example, Hekimoglu et al. [24] quantified uncertainty using an inconsistency score computed as the maximum loss between the initial-refined pairwise task predictions. Simultaneously, a reconstructed feature embedding that condenses information across all tasks measures the diversity score. The combination of these two components forms the basis for deep active learning. Yet this approach cannot handle the domain shift caused by monotask learning from diverse single-label datasets.
@@ III Driving Scene Identification (DSI) Dataset
The data distribution in the high-dimensional scene attribute space is highly imbalanced. Some combinations of scene classes are corner cases difficult to collect data for. This limits the creation of comprehensively annotated datasets to directly train multi-label scene classification models via multitask learning. However, collecting data from respective marginal distributions of scene attributes is relatively straightforward.
This paper introduces DSI, a heterogeneous single-label dataset for demonstrating KAA-CAL, the proposed learning method. DSI consists of 31,835 scene images sampled from public driving video datasets, including BDD100K [7], HSD [8], and ROADWork Data [6], and is supplemented with additional images sampled from YouTube videos. An overview of the DSI dataset and its statistics are presented in Fig. 2. The dataset is available to the public at https://github.com/KELISBU/KAA-CAL.
@@
The DSI dataset consists of $M$ (7) subsets, denoted as $\{\mathcal{D}_{m}\}_{m=1}^{M}$, where $m$ is the index of attributes for scene identification. Each subset is annotated according to one unique scene attribute and further partitioned into training (Trn), validation (Vld), and test (Tst) sets:
$\mathcal{D}_{m}=\mathcal{D}_{m}^{\text{Trn}}\cup\mathcal{D}_{m}^{\text{Vld}}\cup\mathcal{D}_{m}^{\text{Tst}}.$ (1)
@@
Fig. 2 illustrates the scene attributes and respective classes. In total, DSI provides image-level labels for 24 classes, with each falling exclusively under one of the seven scene attributes. Classifying driving scenes according to these scene attributes offers valuable insights for AVs. Notably, DSI can be conveniently expanded by adding new single-label subsets or new classes within existing subsets (e.g., a new design of intersection type) as needs for new knowledge about scene identification emerge.
@@
Grade-separated Infrastructure. Road segments featuring grade-separated infrastructure, such as bridges and tunnels, often present specific challenges for AVs. Bridges above the road may limit the sensor field of view, cause sudden lighting changes, and introduce gusts of wind. Tunnels, on the other hand, have different lighting conditions, limited visibility, and potential congestion. The Grade-separated Infrastructure subset in DSI comprises three classes: Overhead bridges, Tunnels, and Open roads, totaling 7,765 images. This subset was curated from the HSD dataset using query keywords like “overhead bridge” and “tunnel”. To ensure these structures are visually perceivable, additional query criteria such as “Approaching” and “Entering” were used for data sampling.
@@
Road Function. Each road type has specific characteristics, such as a speed limit range, traffic density, vehicle type distribution, terrain, and travel purposes. Thus, understanding the road function assists adaptive driving behavior. The US road functional system has not been systematically annotated in public datasets. To address this, the Road Function subset was created by querying with keywords like “road function” on YouTube and sampling every fifth frame from the collected videos. This dataset consists of 5,740 images across 4 classes: Local, Arterial, Interstate, and Collector.
@@
Weather. Identification of the weather condition is critical because an adverse weather condition risks driving safety and the performance of visual tasks. The Weather subset includes 4,698 frames sampled from the BDD100K dataset, encompassing Overcast, Clear, Foggy, Snowing, and Raining conditions. To mitigate the issue of data imbalance caused by the rarity of foggy scenes, synthetic images were generated to ensure a more balanced distribution across weather conditions.
@@
Work Zone. Driving safety is a concern in work zones, where construction equipment, temporary barriers, road workers, and other related facilities are present. Work zones often involve lane closures. Confusion arises when old, not fully removed lane markings mix with new ones, especially for AVs. Identifying work zones allows AVs to adjust their lane position and speed dynamically, minimizing crash risks. The training and validation data of the Work Zone subset mainly come from BDD100K and HSD, and its testing data are primarily from the ROADWork dataset.
@@
Weather-related Road Condition. The friction coefficient of the road surface changes depending on weather conditions like snow or rain. Identifying these weather-related road conditions allows AVs to adjust their control for better adaptation to the specific surface characteristics. The Weather-related Road Condition subset includes 3,693 images collected from both YouTube and BDD100K, intended for identifying Snowy, Dry, and Wet conditions.
@@
Time of Day. The lighting condition directly impacts the visibility of objects, roads, and the surrounding environment because illumination influences key visual functions like contrast sensitivity, visual acuity, depth perception, and peripheral vision. Consequently, the lighting condition affects not only human drivers’ visual perception and reaction times but also machine vision due to the affected quality of camera-captured data. For example, nighttime and dawn/dusk images often exhibit increased noise, reduced contrast, and color distortion. The Time of Day subset comprises 2,978 images labeled as Night, Daytime, or Dawn/Dusk, sampled from the BDD100K dataset.
@@
Intersection Type. More than half of the combined total of fatal and injury crashes occur at or near intersections due to increased conflict points and complex interactions among vehicles, pedestrians, and other road users. Recognizing the type of intersection allows AVs to anticipate potential conflict points and make informed decisions, such as yielding, stopping, or merging in complex environments. The Intersection Type subset consists of 2,680 images, categorized into four classes: None, 3-way, 4-way, and Roundabout. The data were collected by integrating YouTube videos, which provide roundabout images, along with the HSD and BDD100K datasets, which contribute images of none, 3-way, and 4-way intersections. 3-way and 4-way intersections are typically more noticeable when ego-vehicles are entering the intersection. Therefore, the query keyword “entering” was used when searching for such images in the HSD dataset.